### v1

Abstract
In this study, we investigate the $P$-$V$ criticality and Ruppeiner geometry in the extended phase space of Hayward anti-de Sitter (AdS) black holes. Through thermodynamic analysis, we confirm that Hayward-AdS black holes undergo distinct $P$-$V$ phase transitions and exhibit well-defined critical phenomena in the vicinity of their critical points. These behaviors are characterized by four critical exponents that typically obey the scaling laws predicted by mean-field theory–indicating a consistent thermodynamic framework with classical phase transition systems ($e.g.,$ van der Waals fluids). Furthermore, we employ Ruppeiner geometry to probe the thermodynamic fluctuations of Hayward-AdS black holes, and by calculating the corresponding curvature scalar, we gain direct insights into the interaction nature of the black hole’s microscopic constituents.
I Introduction
Pioneering studies by Hawking and Bekenstein (Bekenstein1972, ; Bardeen1973, ; Hawking:1976de, ) revealed the fundamental nature of black holes—not merely gravitational systems, but entities with distinct thermodynamic properties. Specifically, a black hole’s temperature is determined by its surface gravity, while its entropy is directly associated with the area of its event horizon (Bekenstein1973, ). Crucially, the Hawking temperature of a black hole changes as it accretes matter or emits radiation (Hawking:1974sw, ); this thermodynamic behavior strongly implies that black holes must possess a microscopic structure Strominger:1996sh . Since the prediction of this property, one of the most compelling questions in black hole physics has persisted: what exactly constitutes the microscopic state of a black hole?
Over the past few decades, the field of black hole thermodynamics has witnessed numerous groundbreaking discoveries. Similar to ordinary thermodynamic systems, black holes exhibit a rich variety of phase transitions—among which, those of black holes in anti-de Sitter (AdS) spacetimes have emerged as a focal point of recent research. A core assumption underpinning these studies is the interpretation of the cosmological constant $\Lambda$ as a thermodynamic pressure variable (Kastor:2009wy, ; Dolan:2010ha, ; Kubiznak:2016qmn, ), defined as
$P=-\frac{\Lambda}{8\pi G_{\rm N}}\,,\,\,\,\,\,\,\,\,\,\Lambda=-\frac{(D-1)(D-2)}{2L^{2}}\,$, (1)
where $L$ denotes the radius of the $D$-dimensional AdS spacetime and $G_{\rm N}$ is Newton’s gravitational constant. Within the framework of this definition, a variety of distinct phase transitions emerge in black hole thermodynamics, such as van der Waals Kubiznak:2012wp , (multiple) re-entrant Altamirano:2013ane ; Frassino:2014pha ; Dehyadegari:2017hvd ; Hendi:2017mfu , those with isolated critical points Dolan:2014vba , those with triple points Altamirano:2013uqa , and superfluid-like phase transitions Hennigar:2016xwd . (See Refs.Bhattacharya:2017hfj ; Spallucci:2013osa ; Majhi:2016txt ; Altamirano:2014tva ; Cai:2013qga ; Abdusattar:2023xxs ; Hu:2018qsy for more details on related works and extended analyses of these phase transitions.) [^1] Notably, recent studies have revealed a captivating finding: in black hole systems with quantum conformal anomaly, the scaling laws can be violated under certain conditions Hu:2024ldp ; Abdusattar:2024zzi .
By analyzing the phase transitions of black holes, thermodynamic geometry has emerged as a powerful tool for probing their microscopic structure, with Ruppeiner geometry playing a particularly pivotal role (Ruppeinerb2008, ). The core idea of this approach is to define a line element in the thermodynamic parameter space, which quantifies the “distance” between two adjacent fluctuation states (Ruppeiner95, ). The curvature scalar constructed from this line element directly indicates the nature of interactions among the system’s components: a positive curvature corresponds to repulsive interactions, while a negative curvature signifies attractive interactions. Notably, this interpretation is not unique to black holes, but is derived from the application of Ruppeiner geometry to ordinary thermodynamic systems (Janyszek1990, ; Oshima1999, ), whose cross-system consistency lends reliability to microscopic studies of black holes. Inspired by this, the microscopic origin of black hole thermodynamics has been explored Wei:2012ui ; Wei:2015iwa ; Wei:2019uqg ; Wei:2019yvs ; Guo:2019oad ; Miao:2017fqg ; KordZangeneh:2017lgs ; Wei:2019ctz ; Dehyadegari:2016nkd ; Hu:2020pmr ; Abdusattar:2023xxs . It is important to note that the statistical mechanical study of black holes differs fundamentally from that of ordinary thermodynamic systems: while ordinary systems typically derive macroscopic thermodynamic quantities from the behavior of microscopic particles, black hole research requires “reverse deduction”—inferring information about microscopic structures from known thermodynamic quantities by analyzing macroscopic phase transition behavior.
This paper focuses on the Hayward-AdS black hole, investigating its microstructure through a $P$-$V$ criticality analysis. A well-recognized limitation of classical gravity theories is the existence of spacetime singularities; quantum effects are widely believed to remedy this defect, highlighting the necessity of developing a quantum theory of gravity. However, no mature quantum gravity theory currently exists, making regular black hole models within a semiclassical framework an important avenue for studying singularity correction. Bardeen was the first to propose a singularity-free regular black hole model (bardeen1968non, ), which replaces the singularity of traditional black holes with a de Sitter core. Subsequent studies further demonstrated that the charged version of the Bardeen regular black hole can be derived from Einstein gravity coupled to a nonlinear electrodynamic source (AyonBeato:1998ub, ; AyonBeato:2000zs, ), and can even be interpreted as the gravitational field generated by a nonlinear magnetic monopole. Hayward proposed another representative class of regular black hole solutions (Hayward:2005gi, ): similar to the Bardeen solution, it is a degenerate configuration of the gravitational field of a nonlinear source, carrying a magnetic charge and a free integration constant. In theories with infinite towers of higher-order curvature corrections in dimensions $D\geq 5$, regular black holes with anti-de Sitter asymptotics are constructed Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025ftm ; these black holes, along with related branes and solutions coupled to various electrodynamics, are analyzed regarding regularity, core properties, and thermodynamics, where thermodynamic features linked to regularization parameters are also revealed. Motivated by this work, this paper focuses on analyzing the phase structure of the Hayward-AdS black hole and exploring the repulsive interactions among its microscopic components.
The structure of this paper is organized as follows: Section II provides a brief review of the thermodynamics of static, spherically symmetric Hayward-AdS black holes in an extended phase space. In section III, we will study the critical behavior of the Hayward-AdS black hole, and further discusses the critical exponents associated with $P$-$V$ criticality. Section IV focuses on the thermodynamic curvature scalar of the Hayward-AdS black hole near the critical point, analyzing its divergence behavior. Finally, Section V presents the conclusions and prospects of this study. Throughout the paper, natural units are adopted, with $c=\hbar=k_{B}=1$.
II Thermodynamics of the Hayward-AdS black holes
The action for the Hayward-AdS black hole in dimensions $D\geq 5$ is given by Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025ftm
$I=\int\frac{{\rm d}^{D}x\sqrt{|g|}}{16\pi G_{\rm N}}\bigg[R-2\Lambda+\sum_{n=2}^{\infty}{\alpha}_{n}\mathscr{Z}_{n}\bigg]\,$, (2)
where $g$ denotes the metric determinant, and $R$ is the Ricci scalar curvature, ${\alpha}_{n}$ are arbitrary coupling constants with dimensions of length2(n-1). The densities $\mathscr{Z}_{n}$ are selected by the condition that they possess second-order equations on general spherically symmetric ansätze. Varying the action in Eq.(2) yields the field equations, from which the line element of the static, spherically symmetric Hayward-AdS black hole is given by Hennigar:2025ftm
${\rm d}s^{2}=-f(r){\rm d}t^{2}+{f(r)}^{-1}{{\rm d}r^{2}}+r^{2}{\rm d}\Omega^{2}_{D-2}\,$, (3)
where the metric function $f(r)$ takes the form
$f(r)=1-\frac{r^{2}\mathscr{S}(r)}{1+\alpha\mathscr{S}(r)}\quad\quad\quad\quad\text{with}\quad\quad\mathscr{S}(r)=-\frac{1}{L^{2}}+\frac{m}{r^{D-1}}\,.$ (4)
Here ${\rm d}\Omega^{2}_{D-2}$ represents the line element of a (D2)(D-2)-dimensional sphere, $m$ is an integration constant that is related to the Arnowitt–Deser–Misner (ADM) mass or thermodynamic mass $M$ Bueno:2024eig ; Bueno:2024zsx ; Hennigar:2025ftm
$M$ $=\frac{(D-2)\Omega_{D-2}m}{16\pi G_{\rm N}}\,$, (5)
where $\Omega_{D-2}$ is the volume of a unit sphere $2\pi^{(D-1)/2}/\Gamma\left[(D-1)/2\right]$. The event horizon of this black hole is determined by the condition $f(r_{+})=0$, using which the explicit expression for mass $M$ can be written as Hennigar:2025ftm
$M=\frac{(D-2)\Omega_{D-2}r_{+}^{D-1}}{16\pi G_{\rm N}L^{2}}\left(1+\frac{L^{2}}{r_{+}^{2}-\alpha}\right)\,.$ (6)
The Hawking temperature of the Hayward-AdS black hole on an event horizon is obtained as Hennigar:2025ftm
$T=\frac{1}{4\pi r_{+}^{3}}\left[(D-3)r_{+}^{2}-\alpha(D-1)+\frac{(D-1)(r_{+}^{2}-\alpha)^{2}}{L^{2}}\right]\,$, (7)
and the entropy is
$S=\frac{\Omega_{D-2}r_{+}^{D-2}}{4G_{\rm N}}\,\,{}_{2}F_{1}\!\Big(2,1-\frac{D}{2};2-\frac{D}{2};\frac{\alpha}{r_{+}^{2}}\Big)\,.$ (8)
By interpreting the cosmological constant as a thermodynamic pressure as given in Eq.(1), and its conjugate quantity as a black hole thermodynamic volume
$V=\frac{\Omega_{D-2}r_{+}^{D-1}}{D-1}\,$, (9)
the first law of black hole thermodynamics and the corresponding Smarr formula take the form, respectively Dolan:2014vba ; Hennigar:2025ftm
${\rm d}M$ $=$ $T{\rm d}S+V{\rm dP}+\Psi{\rm d}\alpha\,$, (10)
$(D-3)M$ $=$ $(D-2)TS-2VP+2\alpha\Psi\,.$ (11)
In the above, $\Psi=\left(\partial_{\alpha}M\right)_{S,P}$ is a potential conjugate to the coupling constant $\alpha$. It is worthwhile to mention that according to Eq.(8) and black hole thermodynamic volume formula, the entropy is only a function of area/volume, $i.e.$ $S=S(V)$. This feature of the black hole will be used in the next section.
III $P$-$V$ criticality of Hayward-AdS black hole
In this section, we study the $P$-$V$ phase transition and critical phenomena in the extended phase space of Hayward-AdS black holes. We first investigate the critical points and critical behavior, and then further derive the critical exponents to verify whether these exponents satisfy the scaling laws.
III.1 $P$-$V$ phase transition and critical behavior
By rearranging the expression in Eq. (7) and combining it with Eq. (1), one can derive the thermodynamic equation of state for the $D$-dimensional spherically symmetric Hayward-AdS black hole, which is given by Hennigar:2025ftm
$P=\frac{(D-2)r_{+}^{3}T}{4G_{\rm N}\left(r_{+}^{2}-\alpha\right)^{2}}-\frac{(D-2)\left[(D-3)r_{+}^{2}-\alpha(D-1)\right]}{16\pi G_{\rm N}\left(r_{+}^{2}-\alpha\right)^{2}}\,$, (12)
where $r_{+}=[(D-1)V/\Omega_{D-2}]^{{1}/{(D-1)}}$. In the special case where $\alpha=0$, the equation of state of the Hayward-AdS black hole reduces to
$P=\frac{(D-2)T}{4G_{\rm N}r_{+}}-\frac{(D-3)(D-2)}{16\pi G_{\rm N}r_{+}^{2}}\,$, (13)
which coincides exactly with the equation of state of the Schwarzschild-(A)dS black hole. [^2] Notably, the Schwarzschild-(A)dS black hole exhibits no $P$-$V$ phase transition Dolan:2010ha ; Kubiznak:2016qmn ; Ma:2015llh ; Hansen:2016ayo ; Abdusattar:2022bpg .
The necessary condition for the existence of the $P$-$V$ phase transition is that Kubiznak:2012wp ; Cai:2013qga ; Hu:2018qsy ; Abdusattar:2023pck
$\left(\frac{\partial P}{\partial V}\right)_{T}=\left(\frac{\partial^{2}P}{\partial V^{2}}\right)_{T}=0\,$, (14)
or equivalently
$\left(\frac{\partial P}{\partial r_{+}}\right)_{T}=\left(\frac{\partial^{2}P}{\partial r^{2}_{+}}\right)_{T}=0\,$, (15)
has a critical point solution $T=T_{\rm c},\ P=P_{\rm c},\ V=V_{\rm c}$. This system is known to exhibit a small-large black hole phase transition Hennigar:2025ftm , where the critical point
$V_{\rm c}$ $=$ $\frac{\Omega_{D-2}}{D-1}r_{\rm c}^{D-1}\,,\,\,\,\,\,\,r_{\rm c}=\sqrt{\frac{\Big[3(D-1)+2\sqrt{3}\sqrt{(D-2)D}\Big]\alpha}{D-3}}\,$, (16)
$T_{\rm c}$ $=$ $\frac{(D-3)\left[\sqrt{3}\sqrt{(D-2)D^{5}}+D\left(D(2D-1)+\sqrt{3}\sqrt{(D-2)D}-12\right)-6\sqrt{3}\sqrt{(D-2)D}+9\right]}{2\pi r_{\rm c}\left[2\sqrt{3}\sqrt{(D-2)D^{5}}+3D^{3}-27D-9\sqrt{3}\sqrt{(D-2)D}+27\right]}\,$,
$P_{\rm c}$ $=$ $\frac{\sqrt{3}(D-3)^{2}(D-2)^{3/2}\sqrt{D}}{32\pi G_{\rm N}\alpha\left[D+\sqrt{3}\sqrt{(D-2)D}\right]\left[3D+\sqrt{3}\sqrt{(D-2)D}-6\right]}\,.$
It is evident that the critical points explicitly depend on both the dimensionality $D$ of the black hole and the parameter $\alpha$.
To facilitate the clear visualization of phase diagrams, we first define dimensionless reduced pressure, reduced temperature, and reduced volume as follows
$\widetilde{P}\equiv\frac{P}{P_{c}}\,,\quad\quad\widetilde{V}\equiv\frac{V}{V_{c}}\,,\quad\quad\widetilde{T}\equiv\frac{T}{T_{c}}\,.$ (17)
Substituting these definitions into the equation of state, we obtain the reduced form of the pressure-volume relation
$\widetilde{P}=\frac{(D-2)\left\{\alpha(D-1)+\left(\frac{(D-1)\widetilde{V}V_{\rm c}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}\left[3-D+4\pi\widetilde{T}T_{\rm c}\left(\frac{(D-1)\widetilde{V}V_{\rm c}}{\Omega_{D-2}}\right)^{\frac{1}{D-1}}\right]\right\}}{16\pi G_{\rm N}P_{\rm c}\left[\alpha-\left(\frac{(D-1)\widetilde{V}V_{\rm c}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}\right]^{2}}\,.$ (18)
The corresponding $\widetilde{P}-$$\widetilde{V}$ phase diagram is presented in Fig. 1.
Figure 1: $\widetilde{P}-$$\widetilde{V}$ phase diagram of Hayward-AdS black hole. For $\widetilde{T}>1$, the phase diagram exhibits ideal gas behavior, indicating that the black hole exists in a single unique phase. For $\widetilde{T}<1$, it shows oscillatory behavior, revealing the presence of two distinct phases (small black hole and large black hole). These two phases undergo a first-order phase transition, which terminates at the critical point $(\widetilde{V},\widetilde{P})=(1,1)$. Here we set $\alpha=0.2,G_{\rm N}=1.$
III.2 Critical exponents near the critical point
In the following, we will calculate the critical exponents of the Hayward-AdS black hole that describe the behavior of the thermodynamic quantities near the critical point. In the usual thermodynamic system, there are four critical components, $\widetilde{\alpha}$, $\beta$, $\gamma$, and $\delta$, which are defined as follows Kubiznak:2012wp
$C_{V}\propto|t|^{-\widetilde{\alpha}},~~~~~\Delta_{\widetilde{V}}=\widetilde{V}_{l}-\widetilde{V}_{s}\propto|t|^{\beta},~~~~~\kappa_{T}\propto|t|^{-\gamma},~~~~~\widetilde{P}-1\propto\omega^{\delta},$ (19)
where $t=\widetilde{T}-1$, $\omega=\widetilde{V}-1$. These two reduced parameters characterize the critical behavior of the system in the vicinity of the critical point. Here, $s$ and $l$ correspond to the horizon radii of the small and large black holes, respectively, which are the endpoints of the black hole’s coexistence phase. It is straightforward to verify that these critical exponents satisfy the following thermodynamic scaling laws Kubiznak:2012wp
$\widetilde{\alpha}+2\beta+\gamma=2\,,\quad\quad\quad\quad\quad\quad\quad\widetilde{\alpha}+\beta(1+\delta)=2\,$,
$\gamma(1+\delta)=(2-\widetilde{\alpha})(\delta-1)\,,~~~~~~~~\gamma=\beta(\delta-1)\,.$ (20)
Since the black hole entropy is also determined by its horizon radius, $C_{V}$ (constant-volume heat capacity) vanishes-implying the first critical exponent $\widetilde{\alpha}=0$. To calculate the remaining three critical exponents, it is convenient to expand the equation of state given in Eq. (12) around the critical point
$\widetilde{P}(t,\omega)=1+a_{10}t+a_{11}t\omega+a_{03}\omega^{3}+\mathscr{O}(t\omega^{3},\omega^{4})$ (21)
where the coefficients are
$a_{10}$ $=$ $\frac{(D-2){T_{c}}\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{3}{D-1}}}{4G_{\rm N}{P_{c}}\left\{\alpha-\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{2}{D-1}}\right\}^{2}}\,$, (22)
$a_{11}$ $=$ $-\frac{\left\{(D-2){T_{c}}\left[\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right]^{\frac{3}{D-1}}\left[3\alpha+\left(\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}\right]\right\}}{4\left\{(D-1)G_{\rm N}{P_{c}}\left[\left(\frac{(D-1){V_{c}}}{\Omega_{D-2}}\right)^{\frac{2}{D-1}}-\alpha\right]^{3}\right\}}\,$,\,\,
$a_{03}$ $=$ $-\frac{\left(3D+\sqrt{3}\sqrt{(D-2)D}-6\right)\left(3D+2\sqrt{3}\sqrt{(D-2)D}-3\right)\eta}{\sqrt{3}(D-1)^{3}\sqrt{(D-2)D}\left(D+\sqrt{3}\sqrt{(D-2)D}\right)^{4}\Theta}\,$,\,\,
with
$\eta$ $=$ $\sqrt{3}\left(11\sqrt{(D-2)D^{11}}-13\sqrt{(D-2)D^{9}}-72\sqrt{(D-2)D^{7}}+117\sqrt{(D-2)D^{5}}\right)$
$-D\left(D(D(D(D(41-19D)+114)-333)+189)+27\sqrt{3}\sqrt{(D-2)D}\right)$
$\Theta$ $=$ $\left(2\sqrt{3}\sqrt{(D-2)D^{5}}+3D^{3}-27D-9\sqrt{3}\sqrt{(D-2)D}+27\right)$
all of which are not zero. From Eq.(21), we derive $d\widetilde{P}=(a_{11}t+3a_{03}\omega^{2})d\omega$, with $t$ constant along the isothermal curve. Below the critical temperature ($i.e.,$ $T<T_{c}$), the black hole system exhibits two coexisting phases: the small black hole phase and the large black hole phase. We thus denote the solutions corresponding to these phases as ($\omega_{s}$, $\widetilde{V}_{s}$) (small black hole) and ($\omega_{l}$, $\widetilde{V}_{l}$) (large black hole).
Substituting this into the Maxwell equal-area law Spallucci:2013osa ; Majhi:2016txt ; Bhattacharya:2017hfj ; Lan:2015bia -given by ${P}^{*}\left(\widetilde{V}_{s}-\widetilde{V}_{l}\right)=\int^{s}_{l}\widetilde{P}d\widetilde{V}$, we further obtain
$\int_{\omega_{l}}^{\omega_{s}}\omega d\widetilde{P}=\int_{\omega_{l}}^{\omega_{s}}\omega\frac{d\widetilde{P}}{d\omega}d\omega=\int_{\omega_{l}}^{\omega_{s}}(a_{11}t+3a_{03}\omega^{2})\omega d\omega=0\,$, (23)
where ${P}^{*}$ denotes the reduced pressure corresponding to the straight line connecting the two coexisting phases. Additionally, the endpoints of the small and large black hole phases (analogous to the vapor and liquid phases in classical systems) share the same pressure $i.e.,$ ${P}^{*}=\widetilde{P}_{l}=\widetilde{P}_{s}$ Spallucci:2013osa ; Majhi:2016txt ; Bhattacharya:2017hfj ; Lan:2015bia . This implies
$a_{11}t(\omega_{l}-\omega_{s})+a_{03}(\omega_{l}^{3}-\omega_{s}^{3})=0.$ (24)
Solving Eqs.(23) and (24), yields unique non-trivial solutions (with $\omega_{l}\neq\omega_{s}$) for $t<0$. Using $\widetilde{V}_{s}=1+\omega_{s}$ and $\widetilde{V}_{l}=1+\omega_{l}$, we explicitly derive the analytical expressions for the reduced volumes of the two coexisting phases near the critical point
$\widetilde{V}_{s}$ $=$ $1-\sqrt{-\frac{a_{11}}{a_{03}}t}\,$, (25)
$\widetilde{V}_{l}$ $=$ $1+\sqrt{-\frac{a_{11}}{a_{03}}t}\,$, (26)
where $a_{11}/a_{03}>0$ to ensure that the square root is real. From these expressions, the volume difference between the two phases is the following
$\widetilde{V}_{l}-\widetilde{V}_{s}=2\sqrt{-\frac{a_{11}}{a_{03}}t}\propto|t|^{\frac{1}{2}},$ (27)
which gives the third critical exponent $\beta=1/2$.
The isothermal compressibility $\kappa_{T}$ is calculated as follows
$\kappa_{T}\propto-\left.\left(\frac{\partial\widetilde{P}}{\partial\omega}\right)_{t}^{-1}\right|_{\omega=0}\propto|t|^{-1}\,$, (28)
yielding the critical exponent $\gamma=1$. For the critical isotherm ($t=0$), considering higher-order terms, we have $\widetilde{P}-1=a_{03}\omega^{3}$, which gives the final critical exponent $\delta=3$.
To conclude, we have derived four critical exponents for the Hayward-AdS black hole, which are given by
$\widetilde{\alpha}=0\,,~~~~~~~~\beta=\frac{1}{2}\,,~~~~~~~~~\gamma=1\,,~~~~~~~~~\delta=3\,.$ (29)
In particular, these exponents are identical to those of the van der Waals liquid-gas system and thus satisfy the thermodynamic scaling laws (see Eq.III.2).
IV Thermodynamic curvature via $P$-$V$ criticality
In this section, we focus on the relationship between Ruppeiner geometry (with normalized scalar curvature) and the $P$-$V$ phase transition of the spherically symmetric Hayward-AdS black hole. In thermodynamic information geometry, the rarity of a fluctuation between two thermodynamic states is indicative of their greater separation. Consequently, the line element Wei:2019uqg ; Wei:2019yvs
$dl^{2}$ $=$ $\dfrac{C_{V}}{T^{2}}dT^{2}-\dfrac{\left(\partial_{V}P\right)_{T}}{T}dV^{2}\,$, (30)
contains information regarding the effective interaction that exists between two microscopic fluctuation states. For a specific fluid system, the curvature scalar derived from Eq.(30) serves its microstructure interactions. Notably, the entropy within this information geometry framework, known as Ruppeiner geometry, constitutes the thermodynamic potential.
For the Hayward-AdS black hole, the constant-volume heat capacity $C_{V}=T(\partial_{T}S)_{V}$ equals zero. This is because the entropy $S$ of a black hole is a function of the horizon radius $r_{+}$ $i.e.$, $S(r_{+})$. Consequently, both the line element (30) and the curvature scalar $R$ exhibit divergence. To avoid this divergence, a new normalized curvature scalar $R_{N}$ is defined as Wei:2019uqg ; Wei:2019yvs
$R_{\rm N}=RC_{V}\,.$ (31)
This normalized curvature scalar is more penetrating to investigate the microscopic properties of black holes. After straightforward calculations, the normalized scalar curvature of the Hayward-AdS black hole expressed in terms of reduced parameters as follows
$R_{\rm N}=-\frac{\left[(D-3)\zeta^{2}-\alpha(D+1)\right]\left[\alpha(D+1+12\pi\widetilde{T}{T_{\rm c}}\zeta)+(3-D+4\pi\widetilde{T}{T_{\rm c}}\zeta)\zeta^{2}\right]}{2\left[\alpha(D+1+6\pi\widetilde{T}{T_{\rm c}}\zeta)+\zeta^{2}(3-D+2\pi\widetilde{T}{T_{\rm c}}\zeta)\right]^{2}}$ (32)
with
$\zeta=\left(\frac{(D-1)\widetilde{V}{V_{\rm c}}}{\Omega_{D-2}}\right)^{\frac{1}{D-1}}\,.$
Note that $R_{\rm N}$ explicitly depends on the dimensions of the black hole $D$ and the parameter $\alpha$. We depict the behavior of $R_{\rm N}$ as a function of $\widetilde{T}$ in Fig.2.
Figure 2: Normalized Curvature Scalar $R_{\rm N}$ of Hayward-AdS Black Holes: Behavior along the coexistence phase of small and large black holes for dimensions $D=5,6,7,8$. Here we set $\alpha=0.2,G_{\rm N}=1.$
Figure 2 demonstrates that at the critical point, the normalized thermodynamic curvature $R_{\rm N}$ diverges to negative infinity ($R_{\rm N}\rightarrow$-$\infty$) in both the small and large phases. For $\widetilde{T}_{0}<\widetilde{T}<1$, $R_{\rm N}$ remains negative along the coexistence curve in both the small and large phases, indicating that an attractive interaction dominates. However, along the coexistence curve for the large black hole phase at $\widetilde{T}=\widetilde{T}_{0}$, $R_{\rm N}$ can be positive as the temperature deviates from the critical temperature, implying that the microstructural interaction in the coexistence large phase of black holes undergoes a transition from attractive to repulsive during the phase transition. Additionally, as the temperature deviates from the critical temperature, $\left|R_{\rm N}\right|$ decreases, and most of the parameter space the value of $R_{\rm N}$ is near zero, suggesting that a weak repulsive interaction dominates in Hayward-AdS black holes at low temperatures.
Near the critical point, $R_{\rm N}$ changes dramatically and tends to negative infinity, implying that the black hole microstructure changes rapidly around the temperature $\widetilde{T}_{\rm div}$
$\widetilde{T}_{\rm div}=\frac{\zeta\left[(D-3)\zeta^{2}-\alpha(D+1)\right]}{2\pi T_{\rm c}\left(3\alpha+\zeta^{2}\right)}\,.$ (33)
We also compute the curves where $R_{\rm N}$ changes sign, which has the following simple relation with the above divergent temperature
$\widetilde{T}_{0}=\frac{\widetilde{T}_{\rm div}}{2}\,$, (34)
whose traversal indicates a change between attractive or repulsive interactions of the microstructure.
Figure 3: Hayward-AdS black hole:(a) The divergence curves of $R_{\rm N}$; (b) The sign changing curves of $R_{\rm N}$. Here we set $\alpha=0.2,G_{\rm N}=1.$
From Fig.3(a), it is observed that at the critical point of the $P$-$V$ phase transition, the normalized curvature scalar $R_{\rm N}$ diverges, and the corresponding divergent temperature $\widetilde{T}_{\rm div}$ at this point shows no dependence on the spacetime dimension of the black hole (evidenced by the curves for $D=5,6,7,8$ converging and behaving uniformly near the critical region). Likewise, Fig.3(b) illustrates that the sign-changing temperature $\widetilde{T}_{0}$ at the critical point is also independent of the black hole’s spacetime dimension, as the curves for different dimensions exhibit consistent trends around the critical point.
V Conclusions and Discussion
In this study, we investigate the $P$-$V$ criticality and Ruppeiner geometry of Hayward anti-de Sitter (AdS) black holes, aiming to deepen the understanding of their thermodynamic behavior and microscopic properties. First, through thermodynamic analysis in the extended phase space, we confirm that Hayward-AdS black holes undergo distinct $P$-$V$ phase transitions and exhibit well-defined critical phenomena in the vicinity of their critical points. Notably, the four critical exponents describing these critical behaviors strictly satisfy the scaling laws predicted by mean-field theory, indicating a consistent thermodynamic framework with classical phase transition systems ($e.g.$, van der Waals fluids).
Furthermore, we use Ruppeiner geometry to investigate the microstructures of Hayward-AdS black holes—an approach that also enables probing of their thermodynamic fluctuations. We first calculated the normalized scalar curvature ($R_{\rm N}$), a quantity whose sign encodes information about the dominant microscopic interactions and which is linked to the correlation length near the critical point. Solving for $R_{\rm N}$ allowed us to identify two key temperatures: $\widetilde{T}_{\rm div}$ (the temperature at which $R_{\rm N}$ diverges) and $\widetilde{T}_{0}$ (the temperature at which $R_{\rm N}=0$). Notably, within the coexisting phase volume of small and large black holes, two such divergence temperatures exist, and these temperatures coincide exactly at the critical point. Our analysis confirms the existence of $\widetilde{T}_{0}$ and reveals a clear transition in dominant microscopic interactions: near the critical point, attractive interactions prevail, while repulsive interactions become dominant at lower temperatures. This interaction structure is universal across most AdS black holes, and our results thus provide preliminary but valuable insights into the microstructural properties of higher-dimensional Hayward-AdS black holes—along with direct understanding of the interaction nature of their microscopic constituents.
Acknowledgements
This work is supported by National Natural Science Foundation of China (NSFC) under grant No. 12465012, Kashi University high-level talent research start-up fund project under grant No. 022024002, and Tianchi Talented Young Doctors Program of Xinjiang Uyghur Autonomous Region.